OpenStudy (anonymous):

Use Part 1 of the Fundamental Theorem of Calculus to find the derivative of the function g(x) = 4tan(t)dt integral x to 7

OpenStudy (anonymous):

\[ g(x)=\int \limits_x^7 4\tan(t)\;dt \]We want to find \(g'(x)\)?

OpenStudy (anonymous):

Consider the function \(f(t)\) such that \(f'(t) = 4\tan(t)\). This gives us: \[ g(x) =\int\limits_x^7f'(t)\;dt= f(7)-f(x) \]

OpenStudy (anonymous):

Now looking at \[ g(x)=f(7)-f(x) \]We can differentiate both sides to get: \[ g'(x) = 0-f'(x) \]

OpenStudy (anonymous):

Remember that \(f'(x) = f'(t)\Bigg|^{t=x}= 4\tan(t)\Bigg|^{t=x}= 4\tan(x)\).

OpenStudy (anonymous):

So \[ g'(x) = -4\tan(x) \]